If the hypotenuse of an isosceles right triangle is 2 , what is the area of the triangle?
Answer by swincher4391(1107)

Pythagorean Theorem. Since we have an isosceles triangle, we know that the two sides that are equal are the legs, while the hypotenuse is 2.
a^2 + b^2 = 2^2
a^2 + a^2 = 4
2a^2 = 4
a^2 = 2
a = sqrt(2) <--- base and height
A = (1/2) b * h
A = (1/2) * sqrt(2) * sqrt(2) = 1/2 * 2 = 1